The total cost for producing 1000 units of a commodity is $3.3 million, and the revenue generated by the sale of 1000 units is $5.1 million.

(a) What is the profit on 1000 units of the commodity?
P(1000) = $  million

(b) Assuming  C(q) represents total cost and R(q) represents revenue for the production and sale of q units of a commodity, write an expression for profit
Answer #1

As it has been given that the total cost for producing 1000 units of a commodity is $3.3 million, and the revenue generated by the sale of 1000 units is $5.1 million.

(a) What is the profit on 1000 units of the commodity?
P(1000) =TR-TC

=$5.1 million - $3.3 million

=$1.8 million

(b) Assuming  C(q) represents total cost and R(q) represents revenue for the production and sale of q units of a commodity, write an expression for profit

Profit= R(q)-C(q)

=R(1000)-C(1000)

=$5.1 million - $3.3 million

=$1.8 million